the amount of cloth used to amke four cutains is given by the function A=-4xsquared+40x, where x ia the width of one cutain in feet ad A i sthe total area in square feet. Find the width that maximizes the area of the curtains. What is the maximum area?
Answer by lwsshak3(11628)

the amount of cloth used to amke four cutains is given by the function A=-4xsquared+40x, where x ia the width of one cutain in feet ad A i sthe total area in square feet. Find the width that maximizes the area of the curtains. What is the maximum area?
**
Standard form of equation for parabola that opens downward (parabola has a maximum):
y=-A(x-h)^2+k, (h,k)=(x,y) coordinates of vertex.
..
A=-4x^2+40x
complete the square
A=-4(x^2-10x+25)+100
A=-4(x-5)^2+100
Vertex:(5,100)
ans:
width=5 ft
maximum area=100 sq ft